The charter package^{∗}

Peter Dyballa

1994/04/07

1 Introduction

This file contains the external font information needed to load the Charter font, and a suitable package file to utilize it.

From this file all .fd files (font definition files) for the PostScript fonts with old encoding (OT1) are created, as well as those with the new encoding (T1).

2 Customization

The usual customization will be to add support for additional combinations in this file, such as a derived bold slanted font.

3 The docstrip modules

As always we begin by identifying the latest version of the files in the log file. The explicit spaces are necessary in an .fd file and the \string guards against situations where ` is active.

1 ⟨∗package⟩

 2 \ProvidesPackage{charter}[\filedate\space\fileversion\space
 3 PSNFSS2e LaTeX2e package]
 4 \renewcommand{\rmdefault}{bch}
 5 \def\bfdefault{b}

6 ⟨/package⟩

 7 % \subsection{EC-encoded Bitstream Charter}
 8 %

9 ⟨∗T1⟩

10 \typeout{File\space T1bch.fd\space loading \space Bitstream\space Charter}%
11 \DeclareFontFamily{T1}{bch}{}%
12 \DeclareFontShape{T1}{bch}{m}{n}{<->bchr8t}{}%
13 \DeclareFontShape{T1}{bch}{m}{sc}{<->bchrc8t}{}%
14 \DeclareFontShape{T1}{bch}{m}{it}{<->bchri8t}{}%
15 \DeclareFontShape{T1}{bch}{m}{sl}{<->bchro8t}{}%
16 \DeclareFontShape{T1}{bch}{m}{u}{<->bchru8t}{}%
17 \DeclareFontShape{T1}{bch}{b}{n}{<->bchb8t}{}%
18 \DeclareFontShape{T1}{bch}{b}{it}{<->bchbi8t}{}%
19 \DeclareFontShape{T1}{bch}{b}{sc}{<->bchbc8t}{}%

∗This file has version number 2, last revised 1994/04/07.

20 \DeclareFontShape{T1}{bch}{b}{sl}{<->bchbo8t}{}%
21 \DeclareFontShape{T1}{bch}{b}{u}{<->bchbu8t}{}%
22 \DeclareFontShape{T1}{bch}{bx}{n}{<->bchbre8t}{}%
23 \DeclareFontShape{T1}{bch}{bx}{it}{<->bchbie8t}{}%
24 \DeclareFontShape{T1}{bch}{bx}{sc}{<->bchbce8t}{}%
25 \DeclareFontShape{T1}{bch}{bx}{sl}{<->bchboe8t}{}%
26 \DeclareFontShape{T1}{bch}{bx}{u}{<->bchbue8t}{}%
27 \DeclareFontShape{T1}{bch}{mn}{n}{<->bchrrn8t}{}%
28 \DeclareFontShape{T1}{bch}{mn}{sc}{<->bchrcn8t}{}%
29 \DeclareFontShape{T1}{bch}{mn}{it}{<->bchrin8t}{}%
30 \DeclareFontShape{T1}{bch}{bn}{n}{<->bchbrn8t}{}%
31 \DeclareFontShape{T1}{bch}{bn}{it}{<->bchbin8t}{}%
32 \DeclareFontShape{T1}{bch}{bn}{sc}{<->bchbcn8t}{}%

33 ⟨/T1⟩

34 % \subsection{Old \TeX-encoded Bitstream Charter}

35 % 36 ⟨∗OT1⟩

37 \typeout{File\space OT1bch.fd\space loading \space Bitstream\space Charter}%
38 \DeclareFontFamily{OT1}{bch}{}%
39 \DeclareFontShape{OT1}{bch}{m}{n}{<->bchr7t}{}%
40 \DeclareFontShape{OT1}{bch}{m}{sc}{<->bchrc7t}{}%
41 \DeclareFontShape{OT1}{bch}{m}{it}{<->bchri7t}{}%
42 \DeclareFontShape{OT1}{bch}{m}{sl}{<->bchro7t}{}%
43 \DeclareFontShape{OT1}{bch}{m}{u}{<->bchru7t}{}%
44 \DeclareFontShape{OT1}{bch}{b}{n}{<->bchb7t}{}%
45 \DeclareFontShape{OT1}{bch}{b}{it}{<->bchbi7t}{}%
46 \DeclareFontShape{OT1}{bch}{b}{sc}{<->bchbc7t}{}%
47 \DeclareFontShape{OT1}{bch}{b}{sl}{<->bchbo7t}{}%
48 \DeclareFontShape{OT1}{bch}{b}{u}{<->bchbu7t}{}%
49 \DeclareFontShape{OT1}{bch}{bx}{n}{<->bchbre7t}{}%
50 \DeclareFontShape{OT1}{bch}{bx}{it}{<->bchbie7t}{}%
51 \DeclareFontShape{OT1}{bch}{bx}{sc}{<->bchbce7t}{}%
52 \DeclareFontShape{OT1}{bch}{bx}{sl}{<->bchboe7t}{}%
53 \DeclareFontShape{OT1}{bch}{bx}{u}{<->bchbue7t}{}%
54 \DeclareFontShape{OT1}{bch}{mn}{n}{<->bchrrn7t}{}%
55 \DeclareFontShape{OT1}{bch}{mn}{sc}{<->bchrcn7t}{}%
56 \DeclareFontShape{OT1}{bch}{mn}{it}{<->bchrin7t}{}%
57 \DeclareFontShape{OT1}{bch}{bn}{n}{<->bchbrn7t}{}%
58 \DeclareFontShape{OT1}{bch}{bn}{it}{<->bchbin7t}{}%
59 \DeclareFontShape{OT1}{bch}{bn}{sc}{<->bchbcn7t}{}%

60 ⟨/OT1⟩